5. A sled is pulled by a large work horse. The sled has a mass of 1500kg and the horse pulls with 2500N. If the coefficient of friction is 0.1 determine the acceleration.


1
Expert's answer
2021-11-07T17:08:37-0500

"ma=F-F_{fr}=F-\mu mg,"

"a=\frac Fm-\mu g=0.69~\frac m{s^2}."